How much of a 30% saline solution must be mixed with 30L of a 50%saline solution to produce a 35% saline solution?
Answer by rwm(914)

.3x+.5*30=.35(x+30)
solve for x
x=90
check
.3*90+.5*30=.35(120)
27+15=42
ok